OpenStudy (anonymous):

solve 2/(x+3) - 1/2 = 3/(2(x+3))

OpenStudy (anonymous):

Combine the left hand side of the equation first, what do you get when you do that? \[\frac{ 2 }{ x+3 } - \frac{ 1 }{ 2 } = \frac{ 3 }{2(x+3) }\]

OpenStudy (aonz):

eliminate denominators first so times everything by 2(x+3)

OpenStudy (aonz):

or you can do @domu way ^^

OpenStudy (anonymous):

(4-(x+3))/2(x+3)=3/(2(x+3)) 4-x-3=3 1-x=3 -x=2 x=-2